OpenStudy (anonymous):

Solve x2 + 6x + 7 = 0. x = -1 and x = -5 3 plus or minus square root of 2 negative 3 plus or minus square root of 2 quantity of negative 3 plus or minus square root of 2 all over 2

OpenStudy (wolf1728):

i always use the quadratic formula: x = [-b +-(sqrt(b²-4*a*c))] / 2*a a = 1 b = 6 c = 7 x = [-6 +-(sqrt(36 - 4 * 1 * 7)] / 2 x = [-6 +- sqrt(8)] / 2 x = [-6 + 2 * sqrt(2)] / 2 OR x1 = -3 + sqrt(2) x2 = -3 -sqrt(2) which means the answer is negative 3 plus or minus square root of 2 How's that for an answer? LOL

OpenStudy (anonymous):

lol thanks

OpenStudy (wolf1728):

u r welcome :-)